OpenStudy (anonymous):

A conical tent made of canvas has a base that is 20 feet across and a slant height of 14 feet. To the nearest whole unit, what is the area of the canvas, including the floor? Use 3.14 for pi. A. 754 ft2 B. 1,193 ft2 C. 2,135 ft2 D. 534 ft2

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

If I remember right the area of the conical part is: \[\pi r s\]

OpenStudy (mertsj):

and the area of the base is \[\pi r^2\]

OpenStudy (mertsj):

So calculate those two and add them together fot the total area of the tent.

OpenStudy (anonymous):

Surface Area = pi*10*14 + pi* (10)^2 Surface Area = 140 pi + 100 pi Surface Area = 240 pi Surface Area = 240*3.14 = 753.6 cubic feet approx

OpenStudy (anonymous):

so A

OpenStudy (mertsj):

Look good.

OpenStudy (anonymous):

ok thanks :p

OpenStudy (mertsj):

yw